OpenStudy (anonymous):

solve for x. 0=(10+2x)(16+2x)-320 i keep getting not factorable! help?

OpenStudy (anonymous):

4 x^2+52 x-160= 0 solve using the quadratic formula

OpenStudy (anonymous):

yes i got that too but it doesn't work. trust me. i tried quadratic formula.

terenzreignz (terenzreignz):

It just might not be giving you a 'nice' answer. (Nice here meaning an integer of fraction solution) No matter, a solution is a solution. Now, could you state the quadratic formula, just so we're clear?

OpenStudy (anonymous):

x = -b +/- (square root of:) b^2 + 4ac and everything divided by 2a

terenzreignz (terenzreignz):

b^2 - 4ac in the square-root, not b^2 + 4ac ok? Good, now what are your values for a, b, and c?

OpenStudy (anonymous):

sorry yeah thats what i meant. and values are: a=1, b=13, and c=-40, when factored

terenzreignz (terenzreignz):

Don't think so. Check on Satellite's simplification, and tell me what a, b, and c are.

OpenStudy (anonymous):

0=(10+2x)(16+2x)-320 0=(160+20x+32x+4x^2-320 0=4x^2+52x-160 0=4(x^2+13x-40) 0=x^2+13x-40 do you see where i got the values from?

terenzreignz (terenzreignz):

oh okay then, sorry ^_^ Well then, just plug in :)

OpenStudy (anonymous):

okay :P